Question.1:


An adult has on average 5.4 liters of blood. Assume the variable is normally distributed and has a standard deviation of 0.4. Find the percentage of people who have less than 5.7 litres of blood in their system.



Question.2 :


The average annual salary for all U.S teachers is $ 48,675. Assume that the distribution is normal and the standard deviation is $ 4995. Find the probability that a randomly selected teacher earns:


(i) Between $36000 and $42000 a year.


(ii) More than 40,000 a year.



Question.3:


The average daily jail population in the U.S is 706,242. If the distribution is normal and the standard deviation is 52,145, find the probability that on a randomly selected day the jail population is:


(i) Greater than 750,000


(ii) More than 600,000 but not more than 700,000




1
Expert's answer
2022-05-20T11:48:36-0400

Question1:


"P(X<5.7)=P(Z<\dfrac{5.7-5.4}{0.4})"

"=P(Z<0.75)=0.7734"

Question2:

(i):


"P(36000<X<42000)"

"=P(Z<\dfrac{42000-48675}{4995})"

"-P(Z\le\dfrac{36000-48675}{4995})"

"=P(Z<-1.3363)-P(Z\le-2.5375)"

"\approx0.0851"

(ii):


"P(X>40000)=1-P(Z\le\dfrac{40000-48675}{4995})"

"\approx1-P(Z\le-1.7367)\approx0.9588"

Question3:

(i):


"P(X>750000)=1-P(Z\le\dfrac{750000-706242}{52145})"

"\approx1-P(Z\le0.83916)\approx0.2007"

(ii):


"P(600000<X<700000)"

"=P(Z<\dfrac{700000-706242}{52145})"

"-P(Z\le\dfrac{600000-706242}{52145})"

"=P(Z<-0.1197)-P(Z\le-2.0374)"

"\approx0.4316"